Can you please help show me how to solve this quadratic equation by completing the square.
a^2 - 4a + 5 = 0

Answer by stanbon(75887)

solve this quadratic equation by completing the square.
a^2 - 4a + 5 = 0
---
a^2 - 4a + ? = -5 + ?
a^2 -4a + (-4/2)^2 = -5 + (-4/2)^2
a^2 - 4a + 4 = -5 + 4
a^2 - 4a + 4 = -1
(a-2)^2 = -1
a-2 = i or a-2 = -i
a = 2+i or a = 2-i
======================
Comment: If you have not studied complex numbers the
answer you would get would be "no solution".
======================
Cheers,
Stan H.